Please use engineering economy relations. no excel. Problem 3: A family decided to buy a super ski and water sports boat. They took out an $80,000,5-year, 6% per year, compounde
Answer #1

Solution

Given:

Details of current loan:

Principal = $80,000 ; Time = 5 year ; Interest = 6% p.a,compounded monthly => monthly rate = 0.5%

=> Effective interest rate / effective annual interest rate = [ { [( 0.5 / 100) + 1] ^ (12)} - 1] * (100)

=> 6.16% p.a

It is calculated as : monthly installment = P * r * [ (1+r) ^ n) /(((1+r) ^ n) - 1) ]

where r = annual interest rate / 12 / 100

= 80,000 * (0.00513) * (1.00513 ^ 60) / ((1.00513 ^ 60) - 1)

= 80,000 * (0.00513) * (1.359 / 0.359)

= $1553.575 ~ $1553.57 per month

(b) principal due after 2 months :

principal per month = 80,000 / 60 i.e., $1333.33

So, principal due after 2 months is : 80,000 - ( $1333.33 * 2) => $ 77,333.34

(c) Interest already paid in 1 st 2 installements = (1553.57 - 1333.33) * 2 => $440.48

(d) From month 3 onwards the new monthly installment for new loan repayment is :

Principal = $77,333.34 ; Time = 5 year ; Interest = 4.2% p.a,compounded monthly => monthly rate = 0.35%

=> Effective interest rate / effective annual interest rate = [ { [( 0.35 / 100) + 1] ^ (12)} - 1] * (100)

=> 4.28% p.a

It is calculated as : monthly installment = P * r * [ (1+r) ^ n) /(((1+r) ^ n) - 1) ]

where r = annual interest rate / 12 / 100

= 77333.34 * (0.00356) * (1.00356 ^ 58) / ((1.00356 ^ 58) - 1)

= 77333.34 * (0.00356) * (1.228 / 0.228)

= $1482.79  per month